OpenStudy (anonymous):

A circle has its center at (1, 4) and a radius of 2 units. What is the equation of the circle?

hero (hero):

Hint: Use the equation of a circle formula (h,k) = (1,4) r^2 = 2^2

OpenStudy (anonymous):

(x-1)^2+(y-4)^2=2^2 is the correct equation??

OpenStudy (anonymous):

well it would equal 4 instead of 2 squared. I think.

OpenStudy (anonymous):

You are right.

OpenStudy (anonymous):

Thank you(:

OpenStudy (anonymous):

yw

hero (hero):

2^2 = 4, so that's really the same thing.